OpenStudy (anonymous):

Suppose that the probability a seed will germinate is 85%. What is the probability that 7 of these seeds will germinate when 10 are planted? a. 0.1298 b. 0.385 c. 0.12 d. 0.141

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

So you have to use the Binominal formula \(\sf \huge P(k~in~n~trials)= (^n_k)p^kq^{n-k}\) n = number of trials k = number of successes n – k = number of failures p = probability of success in one trial q = 1 – p = probability of failure in one trial for your question n=trials=10 k=successes=7 n-k=10-7=3 p=85% q=1-0.85=0.15 \(\sf \huge P(7~in~10~trials)= (^{10}_7)(0.85)^7(0.15)^{3}\)

OpenStudy (perl):

got a bit cut off at the end $$\sf \Large P(7~in~10~trials)= (^{10}_7)(0.85)^7(0.15)^{3}$$

OpenStudy (anonymous):

is it 10/7 ?

OpenStudy (anonymous):

0.1298 ?

OpenStudy (aaronq):

@perl thanks @420chiick yes